# Comparison of Extended Lubrication Theories for Stokes Flow

### v2

@@ 2.1 Lubrication Theory
(1) $\frac{\partial p}{\partial x}$ $=\nu\Big(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\Big)-\Big(u\frac{\partial u}{\partial x}+v\frac{\partial u}{\partial y}\Big),$
(2) $\frac{\partial p}{\partial y}$ $=\nu\Big(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\Big)-\Big(u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}\Big),$
@@
Lubrication theory provides a linearized approximation to the Navier-Stokes equations valid for geometries in which the fluid is constrained to a relatively thin domain. For the two-dimensional fluid domain $(x,y)\in[x_{0},x_{f}]\times[0,h(x)]$, the characteristic length scales are $L_{x}=x_{f}-x_{0}$ and $L_{y}=\max h(x)>0$. Given a prescribed constant flux $\mathcal{Q}\neq 0$, the characteristic velocities are $U_{*}=\mathcal{Q}/L_{y}$ and $V_{*}=\mathcal{Q}/L_{x}$, and the characteristic pressure is $P_{*}=\eta U_{*}L_{x}/L_{y}^{2}$. The dimensionless variables are $X=x/L_{x}$, $Y=y/L_{y}$, $H=h/L_{y}$, $U=u/U_{*}$, $V=v/V_{*}$, and $P=p/P_{*}$. The Reynolds number is $\text{Re}=\rho U_{*}{\color[rgb]{0,0,0}L_{x}}/\eta$ and the length scale ratio is $\varepsilon=L_{y}/L_{x}$. The Navier-Stokes equations Eqs. 1 and 2 in dimensionless terms are,
@@ Appendix B The Stokes Equations
The biharmonic formulation of the incompressible Navier-Stokes equations is an effective method of solution for low Reynolds number flows [gupta_new_2005, marner_potential-velocity_2014, sen_4oec_2015, biswas_hoc_2017]. Through introduction of the stream function $\Psi(X,Y)$ satisfying,
@@
the Navier-Stokes equations Eqs. 1 and 2 are expressed as,
@@
When $\text{Re}=0$, the velocity-stream formulation Eq. 35 reduces to the biharmonic Stokes equation $\nabla^{4}\Psi=0$. Note, the characteristic pressure used to formulate Eq. 35 is $\tilde{P_{*}}=\varepsilon^{2}P_{*}=\eta U_{*}/L_{x}$, and thus the Stokes model is not sensitive to surface variation as the lubrication models are.
To compare the Stokes solution with the one dimensional lubrication theory solution, we assume the inlet and outlet flow profiles correspond to a fully developed laminar flow with flux $\mathcal{Q}$. The surface boundary conditions for velocity are Eqs. 6 and 7, and the pressure satisfies Eq. 15. The inlet and outlet velocity profiles are,
@@
The solution $\Psi$, $U$, $V$ to the biharmonic Navier-Stokes equations Eqs. 35 and 34 is determined through an iterative second-order accurate finite difference method as given in [biswas_hoc_2017, gupta_new_2005].
@@
(41) $\frac{\partial P}{\partial X}\Big|_{i,j}=\frac{1}{\Delta X^{2}}\Big(U_{i-1,j}+U_{i+1,j}-4U_{i,j}+U_{i,j-1}+U_{i,j+1}\Big)\\ -\frac{1}{2\Delta X}\Big(U_{i,j}(U_{i+1,j}-U_{i-1,j})+V_{i,j}(U_{i,j+1}-U_{i,j-1})\Big),$

### v3

@@ 2.1 Lubrication Theory
(1) $\frac{\partial p}{\partial x}$ $=\eta\Big(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\Big)-\rho\Big(u\frac{\partial u}{\partial x}+v\frac{\partial u}{\partial y}\Big),$
(2) $\frac{\partial p}{\partial y}$ $=\eta\Big(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\Big)-\rho\Big(u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}\Big),$
@@
Lubrication theory provides a linearized approximation to the Navier-Stokes equations valid for geometries in which the fluid is constrained to a relatively thin domain. For the two-dimensional fluid domain $(x,y)\in[x_{0},x_{f}]\times[0,h(x)]$, the characteristic length scales are $L_{x}=x_{f}-x_{0}$ and $L_{y}=\max h(x)>0$. Given a prescribed constant flux $\mathcal{Q}\neq 0$, the characteristic velocities are $U_{*}=\mathcal{Q}/L_{y}$ and $V_{*}=\mathcal{Q}/L_{x}$, and the characteristic pressure is $P_{*}=\eta U_{*}L_{x}/L_{y}^{2}$. The dimensionless variables are $X=x/L_{x}$, $Y=y/L_{y}$, $H=h/L_{y}$, $U=u/U_{*}$, $V=v/V_{*}$, and $P=p/P_{*}$. The Reynolds number is $\text{Re}=\rho U_{*}L_{x}/\eta$ and the length scale ratio is $\varepsilon=L_{y}/L_{x}$. The Navier-Stokes equations Eqs. 1 and 2 in dimensionless terms are,
@@ Appendix B The Stokes Equations
The biharmonic formulation of the incompressible Navier-Stokes equations is an effective method of solution for low Reynolds number flows [gupta_new_2005, marner_potential-velocity_2014, sen_4oec_2015, biswas_hoc_2017]. Compared with the lubrication approximation, the bi-harmonic Navier-Stokes equations are formulated with a single characteristic velocity $U_{*}$ and a single length scale $L_{x}$. The characteristic pressure is then $P_{*}=\eta U_{*}/L_{x}$. Consequently, the biharmonic Navier-Stokes model is not sensitive to surface variation as the lubrication models are.
With the dimensionless variables $X=x/L_{x}$, $Y=y/L_{x}$, $U=u/U_{*}$, $V=v/U_{*}$ and $P=p/P_{*}$, the Navier-Stokes equations Eqs. 1 and 2 are expressed as
(34) $\frac{\partial P}{\partial X}$ $=\Big(\frac{\partial^{2}U}{\partial X^{2}}+\frac{\partial^{2}U}{\partial Y^{2}}\Big)-\text{Re}\Big(U\frac{\partial U}{\partial X}+V\frac{\partial U}{\partial Y}\Big),$
(35) $\frac{\partial P}{\partial Y}$ $=\Big(\frac{\partial^{2}V}{\partial X^{2}}+\frac{\partial^{2}V}{\partial Y^{2}}\Big)-\text{Re}\Big(U\frac{\partial V}{\partial X}+V\frac{\partial V}{\partial Y}\Big).$
Through introduction of the stream function $\Psi(X,Y)$ satisfying,
@@
the equations Eqs. 34 and 35 are expressed as,
@@
When $\text{Re}=0$, the formulation Eq. 37 reduces to the Stokes equation $\nabla^{4}\Psi=0$.
To solve the bi-harmonic Navier-Stokes equations, we assume the inlet and outlet flow profiles correspond to a fully developed laminar flow with flux $\mathcal{Q}$. The surface boundary conditions for velocity are Eqs. 6 and 7, and the pressure satisfies Eq. 15. The inlet and outlet velocity profiles are,
@@
The solution $\Psi$, $U$, $V$ to the biharmonic Navier-Stokes equations Eqs. 37 and 36 is determined through an iterative second-order accurate finite difference method as given in [biswas_hoc_2017, gupta_new_2005]. The biharmonic equation Eq. 37 is discretized as,
(43) $28\Psi_{i,j}-8\big(\Psi_{i-1,j}+\Psi_{i+1,j}+\Psi_{i,j-1}+\Psi_{i,j+1}\big)\\ +\big(\Psi_{i-1,j-1}+\Psi_{i-1,j+1}+\Psi_{i+1,j-1}+\Psi_{i+1,j+1}\big)\\ =3\Delta X\big(U_{i,j-1}-U_{i,j+1}+V_{i+1,j}-V_{i-1,j}\big)\\ +\frac{1}{2}\Delta X^{2}\text{Re}\Big(V_{i,j}\big(U_{i-i,j}+U_{i+1,j}+U_{i,j-1}+U_{i,j+1}\big)\\ -U_{i,j}\big(V_{i-1,j}+V_{i+1,j}+V_{i,j-1}+V_{i,j+1}\big)\Big)$
The stream-velocity relation (36) is discretised as,
(44) $U_{i,j}$ $=\frac{-3}{4\Delta X}\big(\Psi_{i,j-1}-\Psi_{i,j+1}\big)-\frac{1}{4}\big(U_{i,j-1}+U_{i,j+1}\big)$
(45) $V_{i,j}$ $=\frac{3}{4\Delta X}\big(\Psi_{i-1,j}-\Psi_{i+1,j}\big)-\frac{1}{4}\big(V_{i-1,j}+V_{i+1,j}\big).$
@@
(46) $\frac{\partial P}{\partial X}\Big|_{i,j}=\frac{1}{\Delta X^{2}}\Big(U_{i-1,j}+U_{i+1,j}-4U_{i,j}+U_{i,j-1}+U_{i,j+1}\Big)\\ -\frac{\text{Re}}{2\Delta X}\Big(U_{i,j}(U_{i+1,j}-U_{i-1,j})+V_{i,j}(U_{i,j+1}-U_{i,j-1})\Big),$